OpenStudy (anonymous):

Calculus please help?? Find the equation of the line tangent to f(x) = 3x^2 + 2x + 7 at the point (1, 12).

OpenStudy (mertsj):

@VIbarguen1 Can you find the first derivative?

OpenStudy (anonymous):

f'(x)= 6x+2

OpenStudy (mertsj):

That is the slope function. To find the slope of the tangent at (1,12) replace x with 1. That will be the slope. Then find the equation of the line with that slope which contains the point (1,12). That will be the equation of the tangent at the given point.

OpenStudy (anonymous):

so the slope is 8

OpenStudy (anonymous):

is there a formula I can use to find the equation or is it just guessing?

OpenStudy (mertsj):

Do you remember in algebra 1 when you learned that the equation of a line, in point slope form is: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Replace m with 8, replace x_1 with 1, replace y_1 with 12

OpenStudy (anonymous):

y-12=8(x-1) y-12=8x-8 y=8x-8+12 y=8x+4 is that the answer?

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

thnx soo much